# Not All Directions Matter: Toward Structured and Task-Aware Low-Rank Adaptation

### 15 Mar 2026 v1

@@ Appendix H Analysis of Robustness and Representational Efficiency
Appendix I Fairness of parameter budgets.
Appendix J Layer-wise Gradient Similarity
@@
Parameter-efficient fine-tuning (PEFT) is now a common way to adapt large pre-trained models to downstream tasks. It is especially useful when full fine-tuning is too expensive or hard to deploy. In recent years, PEFT has moved far beyond language models. It is now widely used in vision, multimodal learning, and domain-specific applications Xin et al. (2024); Zhang et al. (2025c); Xiao et al. (2026b); Liu et al. ; Wang et al. (2025a).
The appeal of PEFT is simple. It lets us reuse strong pre-trained backbones while only updating a small set of task-specific parameters. This makes training cheaper and often more stable. In NLP, this idea has been used for classification, reasoning, and instruction tuning. In vision, PEFT has also shown strong value in real downstream settings, especially when data are limited or domains shift. For example, prompt-based adaptation has been used for cross-domain road damage detection, where the model needs to transfer across different real-world environments Zhang et al. (2025e); Xiao et al. (2026a). Similar ideas have also appeared in scientific and biomedical tasks, where the target domain is specialized and labeled data are often scarce. Prior work on drug–target binding affinity prediction shows that prompt-based adaptation can help connect strong backbone representations with domain-specific downstream objectives Zhang et al. (2025d); Xiao et al. (2024); Zhang et al. (2026). This need for lightweight and transferable adaptation is also common in scientific imaging tasks, where the data distribution is often very different from natural images Xiao et al. (2025a). Besides, recent studies further extend PEFT to data-efficient training paradigms Zhao et al. (2025, 2026). From this view, the main question for PEFT is not only how many parameters to update. A more important question is how to update them well. Our method, StructLoRA, takes a step in this direction. Instead of treating all low-rank directions equally, it asks which directions are useful for the task and how updates across layers should work together. We believe this matters even more in hard downstream settings, such as low-resource learning, cross-domain transfer, and multimodal alignment. In these cases, naive low-rank adaptation may introduce noisy updates or inconsistent behavior across layers.

### 25 Apr 2026 v2

@@ Appendix H Analysis of Robustness and Representational Efficiency
Appendix I Fairness of parameter budgets.
Appendix J Layer-wise Gradient Similarity
@@
Parameter-efficient fine-tuning (PEFT) is now a common way to adapt large pre-trained models to downstream tasks. It is especially useful when full fine-tuning is too expensive or hard to deploy. In recent years, PEFT has moved far beyond language models. It is now widely used in vision, multimodal learning, and domain-specific applications Shen et al. (2026); Xin et al. (2024); Zhang et al. (2025d); Xiao et al. (2026b); Liu et al. ; Wang et al. (2025a); Li et al. (2026a, 2025); Zeng et al. (2026); Han et al. (2026); Fang et al. (2025); Zhang et al. (2026b).
The appeal of PEFT is simple. It lets us reuse strong pre-trained backbones while only updating a small set of task-specific parameters. This makes training cheaper and often more stable Dong et al. (2025, 2026); Jiang et al. (2026). In NLP, this idea has been used for classification, reasoning, and instruction tuning Zhang et al. (2025c); Zhao et al. (2025a). In vision, PEFT has also shown strong value in real downstream settings, especially when data are limited or domains shift Yang et al. (2026); Qiu et al. (2026); Zhang et al. (2026a); Hu et al. (2026); Chen et al. (2026, 2025); Li et al. (2026c, b); Ma et al. (2026). For example, prompt-based adaptation has been used for cross-domain road damage detection, where the model needs to transfer across different real-world environments Zhang et al. (2025f); Xiao et al. (2026a). Similar ideas have also appeared in scientific and biomedical tasks, where the target domain is specialized and labeled data are often scarce. Prior work on drug–target binding affinity prediction shows that prompt-based adaptation can help connect strong backbone representations with domain-specific downstream objectives Zhang et al. (2025e); Xiao et al. (2024); Zhang et al. (2026c); Ma et al. (2025a, b, c); Ma (2026). This need for lightweight and transferable adaptation is also common in scientific imaging tasks, where the data distribution is often very different from natural images Xiao et al. (2025a). Besides, recent studies further extend PEFT to data-efficient training paradigms Zhao et al. (2025b, 2026). From this view, the main question for PEFT is not only how many parameters to update. A more important question is how to update them wellZhou et al. (2026). Our method, StructLoRA, takes a step in this direction. Instead of treating all low-rank directions equally, it asks which directions are useful for the task and how updates across layers should work together. We believe this matters even more in hard downstream settings, such as low-resource learning, cross-domain transfer, and multimodal alignment. In these cases, naive low-rank adaptation may introduce noisy updates or inconsistent behavior across layers.